# Topology-induced symmetry breaking: a demonstration in antiferromagnetic magnons on a Möbius strip

### v3

@@ Abstract
Since the total number of sites $N=L/l$ is finite, the magnon eignmodes must be discrete, taking place only at $k=0,\,\pm 2\pi/L,\,\pm 4\pi/L\ \cdots$ as plotted in Fig. 2. Because the spectrum $\omega_{\alpha}(k)\neq\omega_{\alpha}(-k)$ is skewed by $\delta k=\pi/L$ while $\omega_{\beta}(k)=\omega_{\beta}(-k)$ is symmetric, only the $\beta$ modes can form standing waves on the Möbius strip, whereas the $\alpha$ modes cannot. Consider the Néel vector for example, $\delta\bm{n}_{\beta}(k)+\delta\bm{n}_{\beta}(-k)\sim\hat{y}\sin\omega_{\beta}t\cos kx$, where $t$ and $x$ are separated, hence representing a standing wave. For the $\alpha$ branch, $\omega_{\alpha}(-k)=\omega_{\alpha}(k+2\delta k)$, thus time and space do not separate in $\delta\bm{n}_{\alpha}(k)+\delta\bm{n}_{\alpha}(-k)$, prohibiting the formation of standing waves. The disparity between the $\alpha$ and $\beta$ branches is an intrinsic topological property of the AFM magnons on a Möbius strip.
@@
At this point, it is instructive to compare the unique magnon eigenmodes on a Möbius strip with what would become the eigenmodes if the nano-ribbon in Fig. 1(a) is wrapped into a topologically trivial band without twisting (which imposes an ordinary PBC). The latter case only admits the well-known eigenmodes in collinear AFM materials [29] because we have excluded the local curvature effect in our model. To this end, Fig. 2(d) illustrates the right-circular ($R$) and left-circular ($L$) eignmodes, as well as their coherent superposition $L-R$ and $L+R$ featuring elliptical precessions of $\bm{S}_{A}$ and $\bm{S}_{B}$ with opposite chirality, which leads to a linearly-polarized oscillation of the Néel vector $\bm{n}$ without chirality (so does the total spin vector $\bm{S}$). While $L-R$ is locally identical to the $\alpha$ mode shown in Fig. 2(a), it is not accompanied by a built-in $\pi$ phase shift at $k=0$, let alone a spectral shift. The $\beta$ modes are not even locally similar to $L-R$ or $L+R$, which are emergent eigenmodes enabled by the Möbius topology. Unlike $L-R$ and $L+R$, neither $\alpha$ nor $\beta$ can be expressed as linear superposition of $R$ and $L$. This is because a Möbius strip is a non-orientable manifold on which the chirality of spin precessions becomes ambiguous globally, given that $+z$ and $-z$ are indistinguishable. In other words, the eigenspace spanned by $\alpha$ and $\beta$ is not smoothly connected to that spanned by $R$ and $L$; they belong to distinct topological classes.
@@
It is established that in easy-axis AFM materials such as MnF2 [19, 30], $R$ and $L$ are degenerate in energy in the absence of magnetic fields. In easy-plane AFM materials such as NiO, the existence of hard-axis anisotropy breaks the rotational symmetry and lifts the degeneracy, rendering $L-R$ and $L+R$ the magnon eigenmodes [20, 29]. That is to say, symmetry dictates the nature of eigenmodes. The spin Hamiltonian we adopted in this Letter, however, preserves the local rotational symmetry (with respect to the local easy axis). Therefore, the suppression of the circularly-polarized modes, hence the absence of chirality in the Néel vector dynamics, is solely attributed to the Möbius topology. This introduces an intriguing but hitherto unknown mechanism: Non-trivial topology in the real space alone can lead to spontaneous symmetry breaking in the eigenspace of elementary excitations without the aid of symmetry-breaking interactions. The physical consequence manifests as a lifted degeneracy of the eigenmodes which possess a lower symmetry than the Hamiltonian. Because a quantum of angular momentum associated with the $R$-($L$-)circular mode is $+\hbar$ ($-\hbar$), the TISB we found here is followed by the suppression of longitudinal magnon spin currents on a Möbius strip.
To close our discussion, we mention that the lowest $\beta$ mode ($k=0$) is in principle observable via the AFM resonance driven by a microwave, while the $\alpha$ modes are difficult to probe as they are at odds with standing waves. Nonetheless, our findings are amenable to meta-materials such as magnonic crystals, wherein an $\alpha$-mode locally excited by an oscillating magnetic field will propagate in opposite directions but can never form standing waves, leaving a non-reciprocal circulation of energy along the strip which is detectable by optical methods.
@@
Outlook. Even though we have demonstrated the TISB in the AFM magnons on a Möbius strip, the mechanism itself is general and can manifest in other contexts with different quasiparticles or TBCs. For example, we anticipate that the eigenmodes of phonon excitations on a Möbius strip to be linearly polarized while the circularly-polarized chiral phonons are suppressed by the Möbius topology. For photons that are governed by Maxwell’s equations, the axis of circular polarization is parallel to the momentum $\bm{k}$, which defines a distinct geometry compared to ours, calling for a separate investigation. For tight-binding electrons on a Möbius strip, while the Hamiltonian is very similar to ours, the fermionic statistics brings a fundamental distinction in the diagonalization [which does not involve $\sigma_{z}$ as that in Eq. (10)]. Under alternative TBCs such as a Klein bottle, even the AFM magnons could acquire new features of TISB beyond what we showed in this Letter. Therefore, our findings could greatly inspire a broader research endeavor in the near future highlighting the profound impact of real-space topology on the physical nature of elementary excitations.
@@ Acknowledgements.
We thank Qian Niu for helpful discussions. This work is supported by the Air Force Office of Scientific Research under Grant No. FA9550-19-1-0307.

### v4

@@ III Analysis and Discussion
The total number of sites $N=L/l$ is finite, so the magnon eignmodes must be discrete, taking place only at $k=0,\,\pm 2\pi/L,\,\pm 4\pi/L\ \cdots$ as plotted in Fig. 3. Because the spectrum $\omega_{\alpha}(k)\neq\omega_{\alpha}(-k)$ is skewed by $\delta k=\pi/L$ while $\omega_{\beta}(k)=\omega_{\beta}(-k)$ is symmetric, the formation of standing waves by these two branches is quite different. A standing wave features separation of $t$ and $x$ in the wavefunction, which, under the boundary conditions, restricts the way of mode pairing. For the $\alpha$ branch, a pair of magnons of momenta $k_{n\pm}^{\alpha}=\delta k(1\pm n)$ can superimpose and form a standing wave
$\delta\bm{n}_{\alpha}(k_{n+}^{\alpha})+\delta\bm{n}_{\alpha}(k_{n-}^{\alpha})\sim\hat{x}\cos(\omega_{\alpha}t)\cos(n\delta kx),$ (21)
where $n=1,3,5...$ is odd integer representing the number of nodes as shown in Fig. 4(a). For an even integer, it is impossible to separate $t$ and $x$ because of the spectral shift. On the other hand, mode pairing within the $\beta$ branch can only happen between $k_{n\pm}^{\beta}=\pm n\delta k$ with an even integer $n=0,2,4...$ such that
$\delta\bm{n}_{\beta}(k_{n+}^{\beta})+\delta\bm{n}_{\beta}(k_{n-}^{\beta})\sim\hat{y}\sin(\omega_{\beta}t)\cos(n\delta kx),$ (22)
which is shown in Fig. 4(a). Comparatively, the $\beta$ branch bears a similar standing-wave formation as the case of trivial PBC, whereas the $\alpha$ branch is unconventional and unique to the TBC. While the number of nodes in each standing-wave mode is well-defined, their locations indicated by the $\cos(n\delta kx)$ factor in Eqs. (21) and (22) are determined up to a global shift. This is because the Hamiltonian Eq. (5) is translationally invariant (with respect to a relocation of $i=1$).
@@
At this point, it is instructive to compare the unique magnon eigenmodes on a Möbius strip with what would become the eigenmodes if the topologically trivial PBC is imposed on the nano-ribbon in Fig. 2(a). Under the PBC, the nano-ribbon only admits the well-known eigenmodes in collinear AFM materials [19, 20, 21] since we have excluded the local curvature effect from the Hamiltonian. As illustrated in Fig. 3(d), these modes are the right-circular ($R$) and left-circular ($L$) eignmodes, which, by a coherent superposition, can form $L-R$ and $L+R$ featuring elliptical precessions of $\bm{S}_{A}$ and $\bm{S}_{B}$ with opposite chirality, leading to a linearly-polarized oscillation of the Néel vector $\bm{n}$ with null chirality (so is $\bm{S}$). While $L-R$ and $L+R$ are linearly polarized, the eigenspace they span is topologically distinct from that spanned by the $\alpha$ and $\beta$ modes. To be specific, while $L-R$ is locally identical to the $\alpha$ mode, it is not accompanied by a built-in $\pi$ phase shift at $k=0$, let alone a spectral shift. The $\beta$ modes do not even locally resemble $L-R$ or $L+R$; they are emergent eigenmodes engendered by the Möbius topology. Unlike $L-R$ and $L+R$, neither $\alpha$ nor $\beta$ can be expressed as a linear superposition of $R$ and $L$. This is because the Möbius strip is a non-orientable manifold on which the chirality of spin precession becomes ambiguous globally, given that $+z$ and $-z$ are indistinguishable. In other words, the eigenspace spanned by $\alpha$ and $\beta$ is not smoothly connected to that spanned by $R$ and $L$; the two cases feature distinct topological sectors.
The structure of the eigenspace deserves an intuitive explanation. The spin precession on each sublattice can be decomposed into $\ket{\circlearrowright}$ and $\ket{\circlearrowleft}$, while the sublattice degree of freedom is represented by |$\ket{A}$ and |$\ket{B}$. With all constraints relieved, the linear space spanned by the direct products of these vectors is 4D. Under the impact of TBC, however, the magnon eigenspace only covers a 2D subspace. Specifically, the quanta of $\alpha$ and $\beta$ eigenmodes described by Eqs. (17) and (18) can be recast into a suggestive from
[…]
where |$\absolutevalue{u_{\rm small}^{\alpha(\beta)}}<\absolutevalue{u_{\rm large}^{\alpha(\beta)}}$ | with $u_{\rm large}^{\alpha(\beta)}+u_{\rm small}^{\alpha(\beta)}\propto\sqrt{q_{\alpha(\beta)}^{1}}$ and $u_{\rm large}^{\alpha(\beta)}-u_{\rm small}^{\alpha(\beta)}\propto\sqrt{q_{\alpha(\beta)}^{2}}$. Comparatively, the eigenspace under the constraint of ordinary PBC, as illustrated in Fig. 3(d), is spanned by
[…]
which is apparently different from that under the TBC. These two subspaces, in spite of their partial overlap, are distinct and cannot be transformed into each other by a linear combination of the coefficients.
One may wonder what happens when the system size approaches infinity. In this limit, the spectral shift of the $\alpha$ branch vanishes so the two branches become degenerate in the eigenvalues. Nevertheless, the eigenvectors (hence the eigenspace) for each allowed $k$ remain structurally distinct from their counterparts associated with the PBC. In other words, the eigenspaces under TBC and PBC do not merge into each other as the system approaches infinity. In particular, the $\pi$ phase shift of the $\alpha$ mode at $k=0$ is topological protected and independent of the system size.
@@
Now let us inspect the manifestation of the proposed TISB. It is established that in easy-axis AFM materials such as MnF2 [19, 32], $R$ and $L$ are degenerate in energy in the absence of magnetic fields. In easy-plane AFM materials such as NiO, the existence of hard-axis anisotropy explicitly breaks the rotational symmetry in the spin Hamiltonian and lifts the degeneracy, rendering $L-R$ and $L+R$ the magnon eigenmodes [20, 21, 22]. In these common scenarios, the magnon eigenstates exhibit the same symmetry with the spin Hamiltonian, as we only consider the excited states while excluding the ground state that is subject to the spontaneous symmetry breaking and recasts the broken symmetry in the Goldstone modes. By contrast, the Hamiltonian (5) we adopted preserves the local rotational symmetry (with respect to the local easy axis), which is explicitly broken in the magnon eigenstates ascribing to the Möbius TBC. This intriguing phenomenon is a direct manifestation of the TISB, i.e., non-trivial topology in the real space alone leads to lifted degeneracy in the eigenvalues, as well as reduced symmeetry in the eigenspace of elementary excitations, without the aid of symmetry-breaking interactions. Because a quantum of angular momentum associated with the $R$-($L$-)circular mode is $+\hbar$ ($-\hbar$), the TISB we found should be followed by the suppression of longitudinal magnon spin currents on a Möbius strip.
Figure 4: (a) Illustration of the standing wave configurations for the lowest three $\alpha$ ($\beta$) modes with an odd (even) number of nodes $n$. The vertical axis indicates the amplitude of spin oscillation, where a sign difference corresponds to a $\pi$ phase difference. (b) Illustration of the elliptical spin trajectories (on site $i=1$) that smoothly deform with a varying $J_{AF}$. Other sites (not shown) follow a similar pattern of change.
To close this section, we mention that the standing waves formed by the $\alpha$ and $\beta$ modes can be locally excited and distinguished through a microwave. While it is difficult to fabricate and measure an AFM Möbius strip using natural materials, our findings are amenable to engineered meta-materials such as magnonic crystals made of artificial AFM units, which can be as large as centimeters with just hundreds of artificial unit cells. If we align the rf field of a microwave source with the longitudinal and transverse directions of the strip, the $\alpha$ and $\beta$ modes can then be selectively driven with different absorption rates. This is because the rf field directly couples the total spin $\bm{S}$ rather than the Néel vector, and $\bm{S}$ is polarized differently in these two modes [as shown in Fig. 3]. By exciting the lowest few standing-wave modes using a frequency-tunable source, the pattern of nodes shown in Fig. 4(a) should in principle be detectable through time-resolved microscopy [33] and other optical approaches.
IV Exotic Phase
What the $\alpha$ and $\beta$ branches of modes share in common is that the elliptical orbits of spin precessions are always elongated in the $x$-direction (longitudinal direction of the strip). This important property, according to Eq. (18), is attributed to $q_{\alpha(\beta)}^{1}>q_{\alpha(\beta)}^{2}$, where $q_{\alpha(\beta)}^{1}=Q_{\alpha(\beta)}+J_{AF}/2$ and $q_{\alpha(\beta)}^{2}=Q_{\alpha(\beta)}-J_{AF}/2$ with $J_{AF}>0$. On this point, it is instrumental to notice that if $J_{AF}<0$, i.e., the AFM inter-chain exchange coupling turns ferromagnetic, then the opposite relation will be true: $q_{\alpha(\beta)}^{1}<q_{\alpha(\beta)}^{2}$, leading to elongated trajectories in the $y$-direction (transverse direction of the strip). However, an immediate problem with the $J_{AF}<0$ regime is that the AFM ground state may become unstable, which will be reflected as the magnon spectra (eigenvalues) touching zero.
One may naively expect a phase transition when $J_{AF}$ flips sign such that the AFM state yields to the ferromagnetic state and forms a domain wall [23, 24]. However, a closer inspection reveals a quite counter-intuitive result. According to Eqs. (15) and (16), we have
$q_{\alpha}^{1}$ $=K+J_{AF}+J_{F}[1-\cos{(k-\xi\pi/L)l}],$ (25a)
$q_{\alpha}^{2}$ $=K+J_{F}[1-\cos{(k-\xi\pi/L)l}],$ (25b)
but now $J_{AF}<0$ is assumed. At momentum $k=\xi\pi/L$, both $q_{\alpha}^{1}$ and $q_{\alpha}^{2}$ reach minimum, so does the eigenvalue:
$\min[\omega_{\alpha}]$ $=S\sqrt{q_{\alpha}^{1}\left(\frac{\xi\pi}{L}\right)q_{\alpha}^{2}\left(\frac{\xi\pi}{L}\right)}$
$=S\sqrt{K(K+J_{AF})},$ (26)
and similarly, the minimum of $\omega_{\beta}$ occurs at $k=0$, where $\min[\omega_{\beta}]=\min[\omega_{\alpha}]$. One can tell from Eq. (26) that the AFM ground is preserved even for $J_{AF}<0$, as long as $K+J_{AF}>0$. A phase transition must take place at $J_{AF}=-K$, where magnons become infinitely soft and proliferate, melting the AFM configuration.
The surprising region of $-K<J_{AF}<0$ features an exotic phase in which the AFM ground state can stand even though the inter-chain and intra-chain exchange interactions are both ferromagnetic. As schematically illustrated in Fig. 4(b), the elliptical trajectories of sublattice spins deform smoothly while $J_{AF}$ varies from positive to negative, until the threshold $J_{AF}=-K$. An ideal circular polarization for each sublattice is reached at $J_{AF}=0$. In the exotic phase with $-K<J_{AF}<0$, the major axes of spin precessions indeed lie in the $y$-direction. Nonetheless, the magnon eigenspace is still a disparate subspace comparing to that dictated by the PBC, regardless of the sign of $J_{AF}$. While the linear polarization status of the Néel vector and the total spin vector remain the same for both $J_{AF}>0$ and $-K<J_{AF}<0$, the magnitudes of their oscillations are noticeably different in the two phases, which can be pictorially inferred from Fig. 4(b) and be rigorously quantified via the same set of formulas in Sec. II and III.
V Outlook
@@
Even though we have demonstrated the TISB in the AFM magnons on a Möbius strip, the mechanism itself is general and can manifest in other contexts with different quasiparticles or TBCs. For example, we anticipate that the eigenmodes of phonon excitations on a Möbius strip to be linearly polarized while the circularly-polarized chiral phonons are suppressed by the Möbius topology. For photons that are governed by Maxwell’s equations, the axis of circular polarization is parallel to the momentum $\bm{k}$, which defines a distinct geometry compared to ours, calling for a separate investigation. For tight-binding electrons on a Möbius strip, while the Hamiltonian is very similar to ours, the fermionic statistics brings a fundamental distinction in the diagonalization [which does not involve $\sigma_{z}$ as that in Eq. (15)]. Under alternative TBCs such as a Klein bottle, even the AFM magnons could acquire new features of TISB beyond what we showed in this work. Therefore, our findings could greatly inspire a broader research endeavor in the near future highlighting the profound impact of real-space topology on the physical nature of elementary excitations.
@@ Acknowledgements.
We thank Qian Niu for helpful discussions. This work is supported by the Air Force Office of Scientific Research under Grant No. FA9550-19-1-0307 and the UC Regents’ Faculty Development Award.